OpenStudy (superhelp101):

How many electrons in the same atom can share the quantum number n = 2? 2 6 8 10

OpenStudy (superhelp101):

@sidsiddhartha

OpenStudy (superhelp101):

there is this formula but i forgot what it is

OpenStudy (sidsiddhartha):

yeah its=2*(2^n)

OpenStudy (sidsiddhartha):

it will be =2(2^2)=2*4=8

OpenStudy (superhelp101):

yep. thank you soo much!!

OpenStudy (sidsiddhartha):

:)